Find the area of a rectangle if three of its corners are (-0.1,-3.4), (-7.9,-3.4), and (-0.1,6.8).
I know that area is l(w) I also know how to plot these points on a graph. What am I missing here?
Suzie

Answer by lwsshak3(11628)

Find the area of a rectangle if three of its corners are (-0.1,-3.4), (-7.9,-3.4), and (-0.1,6.8).
I know that area is l(w) I also know how to plot these points on a graph. What am I missing here?
***
If you connect the points you plotted, notice that you will get a right triangle.
The vertical leg=6.8+3.4=10.2=length of rectangle
The horizontal leg=7.9-.1=7.8=width of rectangle
Area=length*width=10.2*7.8=79.56 sq units